Can anybody help me with this question? The question is:
Develop a 90 percent confidence interval for the population mean. Interpret the results.
The Greater Pittsburgh Area Chamber of Commerce wants to estimate the mean time workers who are employed in the downtown area spend getting to work. A sample of the 15 workers reveals the following number of minutes spent traveling.
29 38 38 33 38 21 45 34
40 37 37 42 30 29 35
Develop a 98 percent confidence interval for the population mean. Interpret the results.
Thank you.

Answer by stanbon(48568)

Develop a 90 percent confidence interval for the population mean. Interpret the results.
The Greater Pittsburgh Area Chamber of Commerce wants to estimate the mean time workers who are employed in the downtown area spend getting to work. A sample of the 15 workers reveals the following number of minutes spent traveling.
29 38 38 33 38 21 45 34
40 37 37 42 30 29 35
------------
sample mean = 35.067
sample std = 6.017
-------
standard error = z*s/sqrt(n)
standard error = 1.645*6.017/sqrt(15) = 2.556
-------------------------------------------------------
90% CI: 35.067-2.556 < u < 35.067+2.556
90% CI: 32.511 u < 37.62
======================================================
Develop a 98 percent confidence interval for the population mean. Interpret the results.
Use the z-value for 98%: z = 2.3263
sample mean is the same as above
standard error = 2.3263*6.017/sqrt(15)=3.614
---
98% CI: 35.067 - 3.614 < u < 35.067 + 3.614
98% CI: 31.45 < u < 38.68
---------------------------------
Meaning: We have 98% confidence the population mean is between
31.45 and 38.68
==============================================
Cheers,
Stan H.